# Stolz positive scalar curvature structure groups, proper actions and equivariant 2-types

### v2

@@
Abstract
In this note, we study equivariant versions of Stolz’ R-groups, the positive scalar curvature structure groups $R^{spin}_{n}(X)^{G}$, for proper actions of discrete groups $G$. We define the concept of a fundamental groupoid functor for a $G$-space, encapsulating all the fundamental group information of all the fixed point sets and their relations. We construct classifying spaces for fundamental groupoid functors. As a geometric result, we show that Stolz’ equivariant R-group $R^{spin}_{n}(X)^{G}$ depends only on the fundamental groupoid functor of the reference space $X$. The proof covers at the same time in a concise and clear way the classical non-equivariant case.
@@ 1 Introduction
In [Stolz], Stolz organized the classification problem of metrics with positive scalar curvature in a long exact sequence
$\dots\to\mathrm{R}^{spin}_{n+1}(X)\to\mathrm{Pos}^{spin}_{n}(X)\to\Omega^{spin}_{n}(X)\to\mathrm{R}^{spin}_{n}(X)\to\dots$ (1.1)
@@
of cobordism groups, where $X$ is a CW-complex. This long exact sequence includes the well-known spin cobordism $\Omega_{*}^{spin}(X)$; then the structure bordism group of metrics of positive scalar curvature $\mathrm{Pos}_{*}^{spin}(X)$, whose cycles $(f\colon M\to X,g)$ are defined like those of $\Omega_{*}^{spin}(X)$, but adding the geometric secondary structure given by a positive scalar curvature metric $g$ on the smooth compact manifold $M$ with spin structure; finally, the object of study of this note, $\mathrm{R}^{spin}_{*}(X)$ which intuitively arises as a sort of mapping cone construction of the forgetful functor $\mathrm{Pos}^{spin}_{*}(X)\to\Omega^{spin}_{*}(X)$.
@@
An explicit calculation of this last group has not yet been achieved in any case of interest. Nonetheless, Stolz has proved that, if $X=B\pi_{1}(M)$ is the classifying space of the fundamental group of a closed spin manifold $M$ with $\dim(M)\geq 5$, then $\mathrm{R}^{spin}_{\dim M+1}(X)$ acts freely and transitively on the space of concordance classes of positive scalar curvature metrics on $M$. This action is canonical and hence the space of concordance classes inherits canonically (after the choice of a base point) an abelian group structure and a construction of an explicit model for this is given in [XieYuZeidler, Section 4]. This has been used, for instance, in [PSZ, PSZ2, SZ, XieYuZeidler] to give a lower bound on the rank of this affine group and of the moduli space of concordance classes of metrics of positive scalar curvature, where moduli space refers to the quotient by the action of the diffeomorphism group of $M$.
@@
A fundamental step in [SZ] uses the fact, which can be implicitly deduced from [Stolz], that a 2-connected map between CW-complexes, such as the classifying map $u\colon M\to B\pi_{1}(M)$, induces an isomorphism between corresponding $\mathrm{R}^{spin}_{*}$ groups for $\geq 6$.
@@
There is an obvious equivariant reformulation of this story: instead of working with a space $X$ one can work with the a $G$-cover $\bar{X}$ with its free $G$-action. One can then go equivalently go back and forth between structure on $X$ and $G$-invariant structure on $\bar{X}$ (pulling back from $X$ to $\bar{X}$ and quotienting by $G$ from $\bar{X}$ to $X$). This way, for a space $X$ with fundamental group $G$ and universal cover $\widetilde{X}$, one can replace equivalently the Stolz sequence (1.1) by its equivariant version
$\ldots\to\mathrm{R}^{spin}_{n+1}(\tilde{X})^{G}\to\mathrm{Pos}^{spin}_{n}(\tilde{X})^{G}\to\Omega^{spin}_{n}(\tilde{X})^{G}\to\mathrm{R}^{spin}_{n}(\tilde{X})^{G}\to\ldots$ (1.2)
@@
where all the cycles are defined as above, by requiring to come with a free and co-compact and (whenever we have metrics) isometric action of $G$, and all maps are required to be $G$-equivariant. This equivalent reformulation is important to get information about the groups in the Stolz sequence via higher index theory, which has been successfully implemented in [PiazzaSchick] and [Zeidler], compare also [XieYu] and the survey [SchickICM]. The proofs of lower bounds on the rank of the affine group of concordance classes of positive scalar curvature are based on these techniques.
@@
The aim of this note is two-fold. The first main contribution is an explicit and concise proof of the fact that (non-equivariantly) the group $R^{spin}_{*}(X)$ depends for $\geq 6$ only on the fundamental group information of $X$. Secondly, we want to analyze the corresponding statement for the case of general proper $G$-actions. Here, we develop basic tools for this generalization and then study it in the equivariant context of a CW-complex endowed with a proper action of a discrete group. Concretely, for we construct an equivariant version of the Stolz exact sequence, compare Proposition 3.3. One contribution is to give a complete proof of exactness, as we are not aware that this is available in the literature. Our main original result is the fact that the equivariant R-groups do only depend on the equivariant 2-type of the space, compare Theorem 4.1. Along the way, we construct an equivariant analog of the space $B\pi_{1}(X)$, namely a “universal space for a given 2-type, compare Section 5.2.
@@
It should be noted that this relies on surgery constructions requiring enough “room”. This is the reason for the dimension restrictions listed above: the manifolds whose positive scalar curvature metrics are controlled have to be spin manifolds of dimension at least $5$, and the group $R^{spin}_{*}(X)^{G}$ can be treated efficiently if $\geq 6$.
@@ A $G$-CW-complex X is a $G$-space together with a $G$-invariant filtration
$\emptyset=X^{(-1)}\subseteq X^{(0)}\subseteq X^{(1)}\subseteq\dots\subseteq X^{(n)}\subseteq\dots\subseteq\bigcup_{n\geq 0}X^{(n)}=X$
such that $X$ carries the colimit topology with respect to this filtration and for each $n\geq 0$ the space $X^{(n)}$ is obtained from $X^{(n-1)}$ by attaching equivariant $n$-dimensional cells, i.e. there exists a $G$-pushout
[…]
@@
We recall the concept of a family of subgroups which we just have used:
@@
We say that $f$ is cellular if $X$ and $Y$ are $G$-CW-complexes and, denoting by $X^{(k)}$ the k-skeleton of X, one has $f(X^{(k)})\subseteq Y^{(k)}$.
The well known and important cellular approximation theorem extends to the equivariant context (compare [ttd, Theorem 2.1]):
@@
Proposition 2.7.
(compare [ttd, (2.6) Proposition]) Let $f\colon B\to C$ be a $\nu$-connected map between $G$-CW-complexes and $A$ another $G$-CW-complex. Write $A,B]^{G}$ for the set of $G$-homotopy classes of $G$-maps from $A$ to $B$. Then
@@
is surjective (or bijective, resp.) if $\dim A\leq\nu$ (or $\dim A<\nu$, resp.).
@@ 3 The Stolz exact sequence
Let $X$ be a $G$-CW-complex. We then define the following groups.
  • $\Omega^{spin}_{n}(X)^{G}$ is the $G$-equivariant spin bordism group: a cycle here is given by a pair $(M,f)$, where $M$ is an $n$-dimensional spin-manifold with cocompact spin-structure preserving111i.e. with a lift of the action to the Spin principal bundle action of $G$ and with a $G$-equivariant reference map $f\colon M\to X$. Two cycles $(M,f)$ and $(M^{\prime},f^{\prime})$ are equivalent if there is a cocompact spin $G$-bordism $W$ from $M$ to $M^{\prime}$ and there exists a $G$-equivariant reference map $F\colon W\to X$ extending $f$ and $f^{\prime}$.
  • $Pos^{spin}_{n}(X)^{G}$ consists of cycles $(M,f,g)$, where the pair $(M,f)$ is as before and $g$ is a $G$-invariant metric with positive scalar curvature on $M$. Two such cycles $(M,f,g)$ and $(M^{\prime},f^{\prime},g^{\prime})$ are equivalent if there exists a spin bordism $(W,F)$ as before, along with a $G$-invariant metric $g_{W}$ on $W$ which is of product type near the boundary which restricts to $g$ on $M$ and to $g^{\prime}$ on $M^{\prime}$.
  • $R^{spin}_{n}(X)^{G}$ is the bordism group of spin $G$-manifolds with boundary (possibly empty) of dimension $n$, together with a $G$-invariant Riemannian metric of positive scalar curvature on the boundary. Bordisms are then manifolds with corners. In particular, $(M,f,g)$ and $(M^{\prime},f^{\prime},g^{\prime})$ are equivalent if there exists a $G$-bordism $(W,F,\bar{g})$, where $W$ is a bordism between $M$ and $M^{\prime}$ and the resulting bordism $\partial_{0}W$ between $\partial M$ and $\partial M^{\prime}$, carries a $G$-invariant metric $\bar{g}$ with positive scalar curvature, so that it is a bordism between $(\partial M,\partial f,g)$ and $(\partial M^{\prime},\partial f^{\prime},g^{\prime})$ in the sense of $Pos^{spin}_{n-1}(X)^{G}$.
Each of these sets is equipped with an abelian group structure given by disjoint union of manifolds and is covariantly functorial in $X$ as follows: a $G$-equivariant map of $G$-CW-Complexes $\varphi:X\rightarrow Y$ induces a mapping $\varphi_{*}$ on these groups by post-composing the reference maps with $\phi$.
@@
Note that for each cycle $f\colon M\to X$ in $\Omega_{n}^{spin}(X)^{G}$ the isotropy of $M$ is restricted to belong to the family $\mathcal{I}(X)$ (because the image under the equivariant map $f$ of a point $x\in M$ fixed by a subgroup $H$ of $G$ must also be fixed by $H$, and is a point in $X$).
If we want to restrict the isotropy even further, to live in a family $\mathcal{F}$ of subgroups of $G$, we can replace the space $X$ by $X\times E_{\mathcal{F}}G$ where $E_{\mathcal{F}}G$ is the universal $G$-CW-complex with isotropy family $\mathcal{F}$. This space is characterized by the property that $E_{\mathcal{F}}G^{H}$ is empty if $H\notin\mathcal{F}$ and $E_{\mathcal{F}}G^{H}$ is contractible if $H\in\mathcal{F}$, in particular $\mathcal{I}(E_{\mathcal{F}}G)=\mathcal{F}$. It exists for each family $\mathcal{F}$ and it is unique up to $G$-equivariant homotopy equivalence. It has the universal property that a $G$-CW-complex $X$ with isotropy contained in $\mathcal{F}$ has a unique homotopy class of $G$-maps to $E_{\mathcal{F}}G$. These spaces were introduced and studied in [tomDieck, ttd].
@@
$\ldots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${R^{spin}_{n+1}(X)^{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Pos^{spin}_{n}(X)^{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Omega^{spin}_{n}(X)^{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R^{spin}_{n}(X)^{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\ldots}$
where the first map sends a manifold to its boundary, the second one is the forgetful map (i.e. it forgets the metric of positive scalar curvature) and the last one considers a closed manifold as a manifold with empty boundary.
@@
This is a rather direct consequence of the definitions and well known to the experts, at least non-equivariantly, stated e.g. in [Stolz, (4,4)] (but without proof). The argument for the $G$-equivariant case is exactly the same as the classical non-equivariant situation. As we are not aware of a treatment available in the journal literature, we follow the referee’s suggestion and give a complete account of the argument here.
@@
  • Given $W,f,g]\in R^{spin}_{n+1}(X)^{G}$, its image in $\Omega_{n}(X)^{G}$ is represented by $(\partial W,f|_{\partial W})$ which indeed is a boundary, hence represents 0.
  • Given $M,f,g]\in Pos_{n}^{spin}(X)^{G}$, its image in $R^{spin}_{n}(X)^{G}$ is the manifold $M$ with empty boundary. It is bordant in $R^{spin}_{n}(X)^{G}$ to $\emptyset$ and hence represents 0 via the bordism $M\times[0,1]$, where $(M\times\{1\},g)$ is a positive scalar curvature bordism of $\emptyset=\partial M$ to $\emptyset=\partial\emptyset$.
  • Given $M,f]\in\Omega^{spin}_{n}(X)^{G}$, its image in $Pos_{n-1}^{spin}(X)^{G}$ is represented by $\partial M=\emptyset$, hence represents 0.
@@
  • Assume that $M,f]\in\Omega^{spin}_{n}(X)^{G}$ is mapped to 0 in $R^{spin}_{n}(X)^{G}$. This means that there is a null-bordism $(W,F)$. In the case at hand, $\partial W$ has two disjoint parts: on the one hand $(M,f)$ and on the other hand $(M^{\prime},f^{\prime},g^{\prime})$ which is a bordism from $\partial M=\emptyset$ to $\partial\emptyset=\emptyset$ and which is equipped with a metric $g$ of positive scalar curvature. But this means, by definition, that $M,f]=[M^{\prime},f^{\prime}]\in\Omega^{spin}_{n}(X)^{G}$ where $M^{\prime}$ is equipped with a metric $g$ of positive scalar curvature. Hence we have the preimage $M^{\prime},f^{\prime},g]\in Pos^{spin}_{n}(X)^{G}$ of the initial equivariant bordism class $(M,f)$.
  • Assume that $M,f,g]\in Pos^{spin}_{n}(X)^{G}$ is mapped to 0 in $\Omega^{spin}_{n}(X)^{G}$. This means that there is a bordism $(W,F)$ with $\partial(W,F)=(M,f)$. But then $[W,F,g]$ represents a class in $R^{spin}_{n+1}(X)^{G}$ which is mapped to $M,f,g]\in Pos^{spin}_{n}(X)^{G}$.
  • Finally, assume that $W,f,g]\in R^{spin}_{n+1}(X)^{G}$ is mapped to 0 in $Pos_{n}^{spin}(X)^{G}$. This means that there is a second bordism $(W^{\prime},f^{\prime},g^{\prime})$ whose boundary is $(\partial W,f|_{\partial W},g)$. We now consider the closed $G$-manifold $Z:=W\cup_{\partial W}W^{\prime}$ with map $F:=f\cup_{\partial W}f^{\prime}\colon Z\to X$ and the bordism $B:=Z\times[0,1]$ with map $F\circ pr_{Z}\colon B\to X$. The boundary of $B$ consists of three parts: the internal boundary $W^{\prime}$ between $\partial W$ and $\emptyset=\partial Z$, equipped with the metric $g^{\prime}$ of positive scalar curvature extending $g$. The other two boundary parts are $W$ and $Z$, and by definition $(B,F\circ pr_{Z},g^{\prime})$ is a bordism in $R^{spin}_{n}(X)^{G}$ between $(W,f,g)$ and $(Z,f,\emptyset)$ and $Z,f]\in\Omega^{spin}_{n}(X)^{G}$ represents a preimage of $W,f,g]\in R^{spin}_{n}(X)^{G}$.
@@
$\pi\colon(\widetilde{X},\tilde{x}_{0})\to(X,x_{0})$
be the associated universal covering projection. Then the proper $G$-action on $X$ lifts to a $\widetilde{G}-$action on $\widetilde{X}$, where
@@
is an extension of discrete groups defined as follows: the elements of $\widetilde{G}$ are pairs $(\alpha\colon\tilde{X}\to\tilde{X},g)$ where $g\in G$ and $\alpha$ covers the action map of $g$ on $X$. We define the multiplication by composition of the maps $\alpha$ and multiplication in $G$.

### v3

@@
[…]
Stolz Positive Scalar Curvature Structure Groups
[…]
Stolz Positive Scalar Curvature Structure Groups,
Proper Actions and Equivariant 2-Types
[…]
Massimiliano PUGLISI a, Thomas SCHICK b and Vito Felice ZENOBI c
[…]
T. Schick, M. Puglisi and V.F. Zenobi
[…]
Received February 11, 2025, in final form October 19, 2025; Published online October 30, 2025
[…]
In this note, we study equivariant versions of Stolz’ $R$-groups, the positive scalar curvature structure groups $R^{\rm spin}_{n}(X)^{G}$, for proper actions of discrete groups $G$. We define the concept of a fundamental groupoid functor for a $G$-space, encapsulating all the fundamental group information of all the fixed point sets and their relations. We construct classifying spaces for fundamental groupoid functors. As a geometric result, we show that Stolz’ equivariant $R$-group $R^{\rm spin}_{n}(X)^{G}$ depends only on the fundamental groupoid functor of the reference space $X$. The proof covers at the same time in a concise and clear way the classical non-equivariant case.
[…]
positive scalar curvature; universal space for proper actions; spin bordism; fundamental groupoid
[…]
57R91; 57R90; 53C27; 53C21
@@ 1 Introduction
In [10], Stolz organized the classification problem of metrics with positive scalar curvature in a long exact sequence
$\cdots\to R^{\rm spin}_{n+1}(X)\to\mathrm{Pos}^{\rm spin}_{n}(X)\to\Omega^{\rm spin}_{n}(X)\to R^{\rm spin}_{n}(X)\to\cdots$ (1.1)
@@
of cobordism groups, where $X$ is a CW-complex. This long exact sequence includes the well-known spin cobordism $\Omega_{*}^{\rm spin}(X)$; then the structure bordism group of metrics of positive scalar curvature $\mathrm{Pos}_{*}^{\rm spin}(X)$, whose cycles $(f\colon M\to X,g)$ are defined like those of $\Omega_{*}^{\rm spin}(X)$, but adding the geometric secondary structure given by a positive scalar curvature metric $g$ on the smooth compact manifold $M$ with spin structure; finally, the object of study of this note, $R^{\rm spin}_{*}(X)$ which intuitively arises as a sort of mapping cone construction of the forgetful functor
$\mathrm{Pos}^{\rm spin}_{*}(X)\to\Omega^{\rm spin}_{*}(X).$
@@
An explicit calculation of this last group has not yet been achieved in any case of interest. Nonetheless, Stolz has proved that, if $X=B\pi_{1}(M)$ is the classifying space of the fundamental group of a closed spin manifold $M$ with $\dim(M)\geq 5$, then $R^{\rm spin}_{\dim M+1}(X)$ acts freely and transitively on the space of concordance classes of positive scalar curvature metrics on $M$. This action is canonical and hence the space of concordance classes inherits canonically (after the choice of a base point) an abelian group structure and a construction of an explicit model for this is given in [15, Section 4]. This has been used, for instance, in [6, 7, 9, 15] to give a lower bound on the rank of this affine group and of the moduli space of concordance classes of metrics of positive scalar curvature, where moduli space refers to the quotient by the action of the diffeomorphism group of $M$.
@@
A fundamental step in [9] uses the fact, which can be implicitly deduced from [10], that a 2-connected map between CW-complexes, such as the classifying map $u\colon M\to B\pi_{1}(M)$, induces an isomorphism between corresponding $R^{\rm spin}_{*}$ groups for $\geq 6$.
@@
There is an obvious equivariant reformulation of this story: instead of working with a space $X$ one can work with the a $G$-cover $\bar{X}$ with its free $G$-action. One can then go equivalently back and forth between structure on $X$ and $G$-invariant structure on $\bar{X}$ (pulling back from $X$ to $\bar{X}$ and quotienting by $G$ from $\bar{X}$ to $X$). This way, for a space $X$ with fundamental group $G$ and universal cover $\tilde{X}$, one can replace equivalently the Stolz sequence (1.1) by its equivariant version
$\cdots\to R^{\rm spin}_{n+1}\bigl(\tilde{X}\bigr)^{G}\to\mathrm{Pos}^{\rm spin}_{n}\bigl(\tilde{X}\bigr)^{G}\to\Omega^{\rm spin}_{n}\bigl(\tilde{X}\bigr)^{G}\to R^{\rm spin}_{n}\bigl(\tilde{X}\bigr)^{G}\to\cdots,$
@@
where all the cycles are defined as above, by requiring to come with a free and co-compact and (whenever we have metrics) isometric action of $G$, and all maps are required to be $G$-equivariant. This equivalent reformulation is important to get information about the groups in the Stolz sequence via higher index theory, which has been successfully implemented in [5] and [16], compare also [14] and the survey [8]. The proofs of lower bounds on the rank of the affine group of concordance classes of positive scalar curvature are based on these techniques.
@@
The aim of this note is two-fold. The first main contribution is an explicit and concise proof of the fact that (non-equivariantly) the group $R^{\rm spin}_{*}(X)$ depends for $\geq 6$ only on the fundamental group information of $X$. Secondly, we want to analyze the corresponding statement for the case of general proper $G$-actions. Here, we develop basic tools for this generalization and then study it in the equivariant context of a CW-complex endowed with a proper action of a discrete group. Concretely, we construct an equivariant version of the Stolz exact sequence, compare Proposition 3.3. One contribution is to give a complete proof of exactness, as we are not aware that this is available in the literature. Our main original result is the fact that the equivariant $R$-groups do only depend on the equivariant 2-type of the space, compare Theorem 4.1. Along the way, we construct an equivariant analog of the space $B\pi_{1}(X)$, namely a “universal space for a given 2-type”, compare Section 5.2.
@@
It should be noted that this relies on surgery constructions requiring enough “room”. This is the reason for the dimension restrictions listed above: the manifolds whose positive scalar curvature metrics are controlled have to be spin manifolds of dimension at least $5$, and the group $R^{\rm spin}_{*}(X)^{G}$ can be treated efficiently if $\geq 6$.
@@ A $G$-CW-complex X is a $G$-space together with a $G$-invariant filtration
$\varnothing=X^{(-1)}\subseteq X^{(0)}\subseteq X^{(1)}\subseteq\dots\subseteq X^{(n)}\subseteq\dots\subseteq\bigcup_{n\geq 0}X^{(n)}=X$
such that $X$ carries the colimit topology with respect to this filtration and for each $n\geq 0$ the space $X^{(n)}$ is obtained from $X^{(n-1)}$ by attaching equivariant $n$-dimensional cells, i.e., there exists a $G$-pushout
${\displaystyle\bigsqcup_{i\in I_{n}}G/H_{i}\times S^{n-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X^{(n-1)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\displaystyle\bigsqcup_{i\in I_{n}}G/H_{i}\times D^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X^{(n)}.}$
@@
We recall the concept of a family of subgroups which we just have used.
@@
We say that $f$ is cellular if $X$ and $Y$ are $G$-CW-complexes and, denoting by $X^{(k)}$ the k-skeleton of X, one has $f\bigl(X^{(k)}\bigr)\subseteq Y^{(k)}$.
The well-known and important cellular approximation theorem extends to the equivariant context (compare [12, Theorem 2.1]).
@@
Proposition 2.7 (compare [12, Proposition 2.6]).
Let $f\colon B\to C$ be a $\nu$-connected map between $G$-CW-complexes and $A$ another $G$-CW-complex. Write $A,B]^{G}$ for the set of $G$-homotopy classes of $G$-maps from $A$ to $B$. Then
@@
is surjective ((or bijective, respectively)) if $\dim A\leq\nu$ ((or $\dim A<\nu$, respectively)).
@@ 3 The Stolz exact sequence
Let $X$ be a $G$-CW-complex. We then define the following groups:
  • $\Omega^{\rm spin}_{n}(X)^{G}$ is the $G$-equivariant spin bordism group: a cycle here is given by a pair $(M,f)$, where $M$ is an $n$-dimensional spin-manifold with cocompact spin-structure preserving [^1] That is, with a lift of the action to the Spin principal bundle. action of $G$ and with a $G$-equivariant reference map $f\colon M\to X$. Two cycles $(M,f)$ and $\bigl(M^{\prime},f^{\prime}\bigr)$ are equivalent if there is a cocompact spin $G$-bordism $W$ from $M$ to $M^{\prime}$ and there exists a $G$-equivariant reference map $F\colon W\to X$ extending $f$ and $f^{\prime}$.
  • $\mathrm{Pos}^{\rm spin}_{n}(X)^{G}$ consists of cycles $(M,f,g)$, where the pair $(M,f)$ is as before and $g$ is a $G$-invariant metric with positive scalar curvature on $M$. Two such cycles $(M,f,g)$ and $\bigl(M^{\prime},f^{\prime},g^{\prime}\bigr)$ are equivalent if there exists a spin bordism $(W,F)$ as before, along with a $G$-invariant metric $g_{W}$ on $W$ which is of product type near the boundary which restricts to $g$ on $M$ and to $g^{\prime}$ on $M^{\prime}$.
  • $R^{\rm spin}_{n}(X)^{G}$ is the bordism group of spin $G$-manifolds with boundary (possibly empty) of dimension $n$, together with a $G$-invariant Riemannian metric of positive scalar curvature on the boundary. Bordisms are then manifolds with corners. In particular, $(M,f,g)$ and $\bigl(M^{\prime},f^{\prime},g^{\prime}\bigr)$ are equivalent if there exists a $G$-bordism $(W,F,\bar{g})$, where $W$ is a bordism between $M$ and $M^{\prime}$ and the resulting bordism $\partial_{0}W$ between $\partial M$ and $\partial M^{\prime}$ carries a $G$-invariant metric $\bar{g}$ with positive scalar curvature, so that it is a bordism between $\bigl(\partial M,\partial f,g\bigr)$} and $\bigl(\partial M^{\prime},\partial f^{\prime},g^{\prime}\bigr)$ in the sense of $\mathrm{Pos}^{\rm spin}_{n-1}(X)^{G}$.
Each of these sets is equipped with an abelian group structure given by disjoint union of manifolds and is covariantly functorial in $X$ as follows: a $G$-equivariant map of $G$-CW-Complexes $\varphi\colon X\rightarrow Y$} induces a mapping $\varphi_{*}$ on these groups by post-composing the reference maps with $\varphi$.
@@
Note that for each cycle $f\colon M\to X$ in $\Omega_{n}^{\rm spin}(X)^{G}$ the isotropy of $M$ is restricted to belong to the family $\mathcal{I}(X)$ (because the image under the equivariant map $f$ of a point $x\in M$ fixed by a subgroup $H$ of $G$ must also be fixed by $H$, and is a point in $X$).
If we want to restrict the isotropy even further, to live in a family $\mathcal{F}$ of subgroups of $G$, we can replace the space $X$ by $X\times E_{\mathcal{F}}G$ where $E_{\mathcal{F}}G$ is the universal $G$-CW-complex with isotropy family $\mathcal{F}$. This space is characterized by the property that $E_{\mathcal{F}}G^{H}$ is empty if $H\notin\mathcal{F}$ and $E_{\mathcal{F}}G^{H}$ is contractible if $H\in\mathcal{F}$, in particular $\mathcal{I}(E_{\mathcal{F}}G)=\mathcal{F}$. It exists for each family $\mathcal{F}$ and it is unique up to $G$-equivariant homotopy equivalence. It has the universal property that a $G$-CW-complex $X$ with isotropy contained in $\mathcal{F}$ has a unique homotopy class of $G$-maps to $E_{\mathcal{F}}G$. These spaces were introduced and studied in [11, 12].
@@
$\cdots\to R^{\rm spin}_{n+1}(X)^{G}\to\mathrm{Pos}^{\rm spin}_{n}(X)^{G}\to\Omega^{\rm spin}_{n}(X)^{G}\to R^{\rm spin}_{n}(X)^{G}\to\cdots,$
where the first map sends a manifold to its boundary, the second one is the forgetful map ((i.e., it forgets the metric of positive scalar curvature)) and the last one considers a closed manifold as a manifold with empty boundary.
@@
This is a rather direct consequence of the definitions and well known to the experts, at least non-equivariantly, stated, e.g., in [10, long exact sequence (4.4)] (but without proof). The argument for the $G$-equivariant case is exactly the same as the classical non-equivariant situation. As we are not aware of a treatment available in the journal literature, we follow the referee’s suggestion and give a complete account of the argument here.
@@
  • Given $W,f,g]\in R^{\rm spin}_{n+1}(X)^{G}$, its image in $\Omega_{n}(X)^{G}$ is represented by $(\partial W,f|_{\partial W})$ which indeed is a boundary, hence represents 0.
  • Given $M,f,g]\in\mathrm{Pos}_{n}^{\rm spin}(X)^{G}$, its image in $R^{\rm spin}_{n}(X)^{G}$ is the manifold $M$ with empty boundary. It is bordant in $R^{\rm spin}_{n}(X)^{G}$ to $\varnothing$ and hence represents 0 via the bordism $M\times[0,1]$, where $(M\times\{1\},g)$ is a positive scalar curvature bordism of $\varnothing=\partial M$ to $\varnothing=\partial\varnothing$.
  • Given $M,f]\in\Omega^{\rm spin}_{n}(X)^{G}$, its image in $\mathrm{Pos}_{n-1}^{\rm spin}(X)^{G}$ is represented by $\partial M=\varnothing$, hence represents 0.
@@
  • Assume that $M,f]\in\Omega^{\rm spin}_{n}(X)^{G}$ is mapped to 0 in $R^{\rm spin}_{n}(X)^{G}$. This means that there is a null-bordism $(W,F)$. In the case at hand, $\partial W$ has two disjoint parts: on the one hand $(M,f)$ and on the other hand $\bigl(M^{\prime},f^{\prime},g^{\prime}\bigr)$ which is a bordism from $\partial M=\varnothing$ to $\partial\varnothing=\varnothing$ and which is equipped with a metric $g$ of positive scalar curvature. But this means, by definition, that $M,f]=\big[M^{\prime},f^{\prime}\big]\in\Omega^{\rm spin}_{n}(X)^{G}$ where $M^{\prime}$ is equipped with a metric $g$ of positive scalar curvature. Hence we have the preimage $\big[M^{\prime},f^{\prime},g\big]\in\mathrm{Pos}^{\rm spin}_{n}(X)^{G}$ of the initial equivariant bordism class $(M,f)$.
  • Assume that $M,f,g]\in\mathrm{Pos}^{\rm spin}_{n}(X)^{G}$ is mapped to 0 in $\Omega^{\rm spin}_{n}(X)^{G}$. This means that there is a bordism $(W,F)$ with $\partial(W,F)=(M,f)$. But then $[W,F,g]$ represents a class in $R^{\rm spin}_{n+1}(X)^{G}$ which is mapped to $M,f,g]\in\mathrm{Pos}^{\rm spin}_{n}(X)^{G}$.
  • Finally, assume that $W,f,g]\in R^{\rm spin}_{n+1}(X)^{G}$ is mapped to 0 in $\mathrm{Pos}_{n}^{\rm spin}(X)^{G}$. This means that there is a second bordism $\bigl(W^{\prime},f^{\prime},g^{\prime}\bigr)$ whose boundary is $(\partial W,f|_{\partial W},g)$. We now consider the closed $G$-manifold $Z:=W\cup_{\partial W}W^{\prime}$ with map $F:=f\cup_{\partial W}f^{\prime}\colon Z\to X$ and the bordism $B:=Z\times[0,1]$ with map $F\circ\operatorname{pr}_{Z}\colon B\to X$. The boundary of $B$ consists of three parts: the internal boundary $W^{\prime}$ between $\partial W$ and $\varnothing=\partial Z$, equipped with the metric $g^{\prime}$ of positive scalar curvature extending $g$. The other two boundary parts are $W$ and $Z$, and by definition $(B,F\circ\operatorname{pr}_{Z},g^{\prime})$ is a bordism in $R^{\rm spin}_{n}(X)^{G}$ between $(W,f,g)$ and $(Z,f,\varnothing)$ and $Z,f]\in\Omega^{\rm spin}_{n}(X)^{G}$} represents a preimage of $W,f,g]\in R^{\rm spin}_{n}(X)^{G}$.
@@
$\pi\colon\ \bigl(\tilde{X},\tilde{x}_{0}\bigr)\to(X,x_{0})$
be the associated universal covering projection. Then the proper $G$-action on $X$ lifts to a $\widetilde{G}-$action on $\tilde{X}$, where
@@
is an extension of discrete groups defined as follows: the elements of $\widetilde{G}$ are pairs $\bigl(\alpha\colon\tilde{X}\to\tilde{X},g\bigr)$ where $g\in G$ and $\alpha$ covers the action map of $g$ on $X$. We define the multiplication by composition of the maps $\alpha$ and multiplication in $G$.